OpenStudy (anonymous):

How do you solve 2650= -x^2+105x?

OpenStudy (hba):

Umm, x^2-105x+2650=0 Here a=1,b=-105 and c=2650 when we compare it to standard equation of quad formula :) \[ax^2+bx+c=0\] Now use Quadratic formula, \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] Just plug in the values and find x :)

OpenStudy (anonymous):

how would I do that?

OpenStudy (hba):

x^2-(SOR)x+POR=0

OpenStudy (anonymous):

I never learned the quadratic formula and how to use it., how do you do that?

OpenStudy (hba):

I gave you the formula You just put the values a=1,b=-105 and c=2650 in it and solve for x :)

OpenStudy (anonymous):

but when I put in the values it wanted the square root of 425, its not a perfect square, what do I do?

OpenStudy (hba):

Leave it as it is.

OpenStudy (anonymous):

But it wants me to fill in a blank for an answer and it says it should be just one number,

OpenStudy (hba):

So you have options ?

OpenStudy (anonymous):

I don't think that's the right equation. The original problem is: A company's monthly profit, P, from a product is given by P = −x2 + 105x − 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550?

OpenStudy (anonymous):

No options

OpenStudy (hba):

I don't want to live on this planet anymore :P

OpenStudy (anonymous):

I started with that then I added 1050 to both sides and then the equation was 2650=-x^2+105x?

OpenStudy (hba):

Now a=1,b=-105 and c=2600

OpenStudy (hba):

1550+1050=2600

OpenStudy (hba):

@hero lol :P

OpenStudy (hba):

@Hero You are playing two methods here not me,so you are confusing her :P

OpenStudy (anonymous):

I'm confused???

OpenStudy (hba):

Haha lol :P @hero See what you did ^^

OpenStudy (anonymous):

How am I supposed to solve this problem ???

OpenStudy (hba):

a=1,b=-105 and c=2600 in Quadratic formula

OpenStudy (anonymous):

I tried the quadratic formula, idk how this relates to the original question, I think it wants one number for an answer. The question was: A company's monthly profit, P, from a product is given by P = −x2 + 105x − 1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1,550?

OpenStudy (hba):

@noelleallheart11 Now your values have changed try it again

OpenStudy (anonymous):

But it won't let me enter a square root as part of the answer, and 425 is not a perfect square

OpenStudy (anonymous):

How am I supposed to find the original question so that its just one number that isn't a fraction or decimal for an answer?

OpenStudy (hba):

She used c= 2650 the first time @Hero when she solved it.

OpenStudy (anonymous):

Isn't that what it's supposed to be?

OpenStudy (hba):

No c=2600

OpenStudy (hba):

So now try a=1,b=-105 and c=2600 in Quad.Formula

OpenStudy (anonymous):

whats the quadratic formula again?

OpenStudy (hba):

Scroll up please

OpenStudy (hba):

You can use breaking the middle term as shown by @Hero